OpenStudy (ashchu117):

@HelpOfTheGods KACAW KACAW!!!!!! Solve for x: 2 over 5 (x - 4) = 2x. A. -8 B. negative 1 over 2 C. 1 D. -1

OpenStudy (anonymous):

The gods have heard your calls

OpenStudy (ashchu117):

XD

OpenStudy (anonymous):

lol i like the way u put kacaw xD anyways lets start

OpenStudy (ashchu117):

ok lol

OpenStudy (anonymous):

is this 2over 5 meaning the fraction 2/5?

OpenStudy (anonymous):

ya thats what im guessing

OpenStudy (ashchu117):

yes @yomamabf

OpenStudy (anonymous):

im thinking maybe c but im not sure

OpenStudy (ashchu117):

ok ill wait

OpenStudy (anonymous):

well it was just a guess based on what the question is asking

OpenStudy (anonymous):

i got -2

OpenStudy (anonymous):

*the Gods Have SPOKEN*

OpenStudy (anonymous):

kacaw me if u need anything else :D

OpenStudy (anonymous):

haha this made me crack up a little, anyways 2/5 (x - 4) = 2x why dont we first start off breaking down the multiple answears. Well, lets eliminate B because we know its not that for sure. Were left with A,C,D, We know its going to be a positive answear due to you cant subtract two negatives. You need a positive and a negative. So what does that give us? eliminate A,B,D and all we have is C. So, C is the answear to your queston. Hope this helped out mate. Good luck

OpenStudy (ashchu117):

lol ok XD bye @HelpOfTheGods

OpenStudy (ashchu117):

thanks @-StudyShark- !

OpenStudy (here_to_help15):

2/5 (x-4) = 2x multiply both sides by 5 to get rid of the fraction 2(x-4) = 10 x distribute 2x -8 = 10 x -8 = 8 x x = -1 check 2/5(-5) =? -2 -2 = -2

OpenStudy (ashchu117):

@Here_to_Help15 um.....

OpenStudy (here_to_help15):

yes?

OpenStudy (ashchu117):

i didn't understand like half of that :(

OpenStudy (here_to_help15):

did i do something wrong?? ill check it again

OpenStudy (amistre64):

2/5 (1-4) = 2(1) not quite

OpenStudy (here_to_help15):

wait did i do tis right or wrong i checked it???

OpenStudy (ashchu117):

man am i confused .-.

OpenStudy (amistre64):

study shark babbled off some inane jibberish ...

OpenStudy (here_to_help15):

i got D?

OpenStudy (here_to_help15):

lol

OpenStudy (amistre64):

here steps are correct. but we would prefer more of the askers participation and more of the responders guiding towards a solution

OpenStudy (ashchu117):

ok o.o

OpenStudy (ashchu117):

im still lost as to what the answer is. plz explain @amistre64

OpenStudy (anonymous):

yea me too can u go through it step by step

OpenStudy (amistre64):

here-to-help did a nice step by step. they should be the one to walk you thru it since they seem to have a good handle on a process.

OpenStudy (ashchu117):

@Here_to_Help15 plz come and explain

OpenStudy (amistre64):

its just using basic algebra properties multiply thru by 5 to undo the fraction divide by 2 to get rid of that 2 then the rest is addition/subtraction with a division to clear it out

OpenStudy (here_to_help15):

its d

OpenStudy (ashchu117):

ok thanks @Here_to_Help15

OpenStudy (here_to_help15):

yw